Nuclear Target Test

From clas12-run
Jump to navigation Jump to search
Role Phone Number
Hall B Run Coordinator (757) 575-7540 (cell)
Hall B Physics Division Liaison (757) 876-1789 (cell)
MCC 7048
Crew Chief 7045
Crew Chief (757) 876-3367 (cell)
Program Deputy (757) 876-7997 (cell)
RadCon (757) 876-1743 (cell)
Gate House Guard 5822
Location Phone Number
Hall B Floor 5165
Hall B Space Frame 5170 and 5171
Hall B Forward Carriage 5371
Hall B Gas Shed 7115
Hall B Counting House 5244 (Shift Expert)
Hall B Counting House (757) 329-4846 (Shift Expert cell)
Hall B Counting House 5245 or 5126 (Shift Worker)
Hall A Counting House 5501
Hall C Counting House 6000
Hall D Counting House 5504
Hall B System Phone Number On-Call Person
Engineering (757) 748-5048 (cell)
(757) 897-9060 (cell)
Engineering On-Call (primary)
Denny Insley (secondary)
Slow Controls (757) 748-6922 (cell) Nathan Baltzell
Beamline (757) 303-3996 (cell) Eugene Pasyuk
DC (757) 218-4372 (cell)
(757) 748-5048 (cell)
Florian Hauenstein (primary)
Engineering On-Call (secondary)
SVT/MVT/MM (757) 541-7539 (cell)
(757) 753-7769 (cell)
Yuri Gotra (primary)
Rafo Paremuzyan (secondary)
ECAL (757) 810-1489 (cell) Cole Smith
FTOF/CTOF (757) 344-7204 (cell) Daniel Carman
CND (757) 329-4844 (cell) Pierre Chatagnon
HTCC/LTCC (757) 344-7174 (cell) Youri Sharabian
FT (757) 344-1848 (cell) Raffaella De Vita
BAND (757) 310-7198 (cell) Florian Hauenstein
RICH (757) 344-3235 (cell)
(757) 748-6922 (cell)
Valery Kubarovsky (primary)
Nathan Baltzell (secondary)
DAQ (757) 232-6221 (cell) Sergey Boiarinov
Cryotarget (757) 218-2266 (cell)
(757) 871-5374 (cell)
(757) 897-1415 (cell)
(757) 746-9277 (cell)
Xiangdong Wei (first)
James Brock (second)
James Maxwell (third)
Chris Keith (fourth)
RG-E Solid Target (757) 344-1848 (cell) Antonio Radic
Authorized Hall B Solenoid/Torus Operators
Hall B Denny Insley, Morgan Cook, Eugene Pasyuk
Magnet Group Probir Ghosha, Renuka Rajput-Ghoshal
Detector Support Group Brian Eng, Pablo Campero, Tyler Lemon
DC Power Onish Kumar, Sarin Philip
  • Note, all non-JLab numbers must be dialed with an area code. When calling from a counting-house landline, dial "9" first.
  • To call JLab phones from outside the lab, all 4-digit numbers must be preceded by 757-269
  • Click Here to edit Phone Numbers. Note, you then also have to edit the current page to force a refresh.

Click Here to edit Phone Numbers. Note, you then also have to edit this page to force a refresh.

CLAS12 Nuclear Target Test, January 31 - February 7, 2020
Beam energy ~10.4 GeV (5 pass)
Important: Document all your work in the logbook!

RC: Eugene Pasyuk

PDL: Maurizio Ungaro

CLAS12 configuration

  • Torus at full field, electrons inbending
  • Solenoid at full field.
  • BMT is removed
  • FTT is OFF
  • SVT is removed and replaced with two sensors
  • BAND is removed
  • FT is not in DAQ configuration. We will use scalers only

Run Plan: Feb 4-5, 2020

  • target is 125um Pb
  • Establish beam to the tagger
  • Establish beam to the Faraday cup
  • Turn all CLAS detectors ON
  • Check beam centerin on FTCAL scalers. Adjust if needed.
  • Start DAQ: PROD66_NUCL_TEST, trigger: rgf_inbending_v1_1.trg
  • Start luminosity scan from 5 nA up to 100 nA with 5 nA steps keep DAQ running
  • For each beam current:
    • reset monitoring histograms
    • make snapshots of:
      1. beam overview
      2. beam strip charts
      3. DAQ status
      4. Run Control
      5. trigger
      6. HV overview
      7. FT scalers
      8. RICH scalers
      9. SVT currents
      10. DC currents
      11. DC HV map
      12. FTOF currents (FTOF Ammeter camera)
    • after 5 minutes of running at set current send monitoring histograms to the logbook
  • some detectors may trip during luminosity scan. Make a note what tripped and at what current. Don't reset, keep going
  • Once you get to 100 nA call RC for further guidance.
  • At 6 am Bob will come to install dosimeters
  • request escorted controlled access to install dosimeters and foils.
  • take beam until noon at TBD nA. Take snapshots of beam overview, beam strip charts, SVT currents strip chart every hour.
  • Bob and RadCOn people are in the counting room request controlled access with RadCon to take dosimeters out of the hall and do activation counting
  • ramp solenoid down and go to restricted access

  • Please note that we have added two items to the list of tasks for each run.
    • (1) When you collect the gui snapshots for the logbook (scalar gui, FTC Asym plot, livetime, ROC rates, beamline stripcharts) include the DAQ gui AND make sure the snapshot is taken when you're running. We want to record the rate plots on the DAQ gui.
    • (2) Measure the trigger structure (trigger-time distribution) about every 30 minutes and at least once per run enter the results in the logbook. See below for instructions. This is needed to monitor the injector laser phase which created serious problems last week.
  • If the trigger-time distribution or region 1 DC occupancies get worse ( > 6.5%) call the RC. What we will likely do is reduce the beam and restart.
  • At some later time we need to do an empty target run. Current will be 150 nA for 50M triggers which should take 2-3 hours.


Trigger file : rtrigger: rgf_inbending_v1_1.trg

Run database: when starting/stopping a run, enter (or check if already present) the requested current in the pop-up window and put RELEVANT COMMENTS (eg: "production run", "trigger test vXX, changed this and that" not just "run") -- please label junk runs if you get the chance. Please also keep a record of run numbers and their conditions in the shift summary. The information you enter in the dialogue box goes directly into the run database -- it's our primary source of run info and mistakes are hard to correct a posteriori.

Please run the script to look at trigger structure once per shift and check that you get a smooth distribution. Instructions are here: [1]

Keep an eye on CND channel S14L1R -- it's a bit temperamental. If it trips, switch HV off and on again for it and make an entry in the log book.

At the end of each run, follow the STANDARD DAQ RESTART SEQUENCE
This needs to be followed at the end of every run. Note: this might be different from the instructions coming up on RunControl.

  • "end run", "cancel", "reset"
  • if the run ended correctly:
    • "download", "prestart", "go"
  • if the run did not end correctly or if any ROCs had to be rebooted:
    • "configure", "download", "prestart", "go"
  • After each step, make sure it completed in the Run Control message window. If a ROC has crashed, find which one it is, issue a roc_reboot command ON JUST THAT ROC and try again. Contact the DAQ expert if there are any questions.

References and Standards:

(last update 1/29/2020- 23:20)

Nominal Beam Positions

  • 2H01, X: -1.0 mm, Y: +1.5 mm
  • 2C21, X: 0.0 mm, Y: 0.1 mm
  • 2C24, X: -0.3 mm, Y: +0.2 mm

FSD Thresholds

  • Upstream: 2000 Hz
  • Midstream: 2000 Hz
  • Downstream: 1500000 Hz
  • BOM: 500000
  • Integration time: 5ms

Reference Harp Scans for Beam on Faraday Cup: 2C21 [2], 2C24 (tagger harp) [3], 2H01 [4]

Reference Harp Scans for Beam on Tagger dump: 2C21 [5], 2C24 (tagger harp) [6]

Reference Monitoring Histograms (40 nA, production trigger) [7]

Counter rates

  • Upstream counters integrated rates: 0-15 Hz (acceptable up to 100 Hz are acceptable) @50 nA.
  • Midstream counters: 10-20 Hz (acceptable up to 50 Hz) @50 nA.
  • Counting rates of ~ hundreds of Hz may indicate bad beam tune or bleed-through from other Halls.

Beamline vacuum: should be not higher than 5e-5.

General Instructions:

  • The main lights in the Hall (dome lights) and the Forward Carriage lights are being kept off because of light leaks affecting some of the detectors. If these lights are switched on during an access, they should be switched off when leaving the Hall. Note that the dome lights when switched off cannot be turned back on immediately because they require 10-15 min to cooldown.
  • Do not run more than 30 minutes above 30 nA with 5-pass beam without the beam blocker in front of Faraday cup . Put beam blocker in for long running at high currents for 5 pass operations.
  • Turn DC HV off only for beam tuning; if no beam is available or when beam is stable, keep them on even if you are not taking data.
  • In case of loss of communication with IOCBTARG , follow instructions at
  • With any issue contact On-Call Experts or RC - do not spend more than 15-20 min trying to fix the problem. If the phone is not answered, try the "Page Experts" button on top of CLASCSS.
  • Check for and read any comments to log-book entries made during your shift!

Every Shift:

  • Follow run plan as outlined by RC
  • If any concern about beam stability, ask MCC if orbit locks are on (they should be).
  • Keep shift summary up to date in HBLOG. Record all that happens.
  • Document any beam condition change and send scaler GUIs to HBLOG
  • Fill out BTA hourly. Click "Load from EPICS" to automatically fill the left side. Run the script to print for you the correct BTA info for this hour -- edit the values on the BTA page if they are wrong!
  • Upload spectra from clas12mon at begin and end of shift and cross-check with reference spectra (make sure you select the SAME SECTOR as for the reference plots on tabs which have multiple sectors).
    • Reset clas12mon every ~2 hours and inspect spectra (some detectors have buttons at the bottom to toggle between the sectors -- please check all of them).
  • Fill and submit the shift checklist in the logbook
  • Run script to check smoothness of the trigger time structure: [8]
  • Monitor the SVT Slow Controls status, post plot of current stability in the HBSVT elog once per shift.

Every Run:

  • Check values at the start of every run and log screenshots of:
    • Main scaler GUI display
    • Trigger rate GUI
    • Beam strip charts
    • FTC scalers
  • At the start and end of every run, check that the data filled in the database run list is correct (especially fields which are filled in manually, like beam current and comments). This replaces the previously-used spreadsheet.
  • When checking the ECAL response, make sure to look at the histograms for each sector by using the sector buttons at the bottom of the clas12mon window - if adcecal* or/and adcpcal* ROC crashes, there will be no alarm and the histograms are the way to notice this issue.
  • Same holds for FTOF.

To access the Hall:

  • If you need to go down into Hall B alert Denny Insley (757-897-9060, cell/text: 757-810-9576) and the Run Coordinator (757-575-7540).
  • Check in with the shift personnel before entering the hall.

Notes / Troubleshooting :

If crate ADCBAND was rebooted: Contact BAND expert and make a note in the HBBAND logbook and in HBLOG.

FTOF HVs: The goal is to minimize the number of power cycles of the dividers.

    • OFF during: initial beam tuning to the Faraday Cup after CLAS12 has been off for a long shutdown, during a Moller run, harp scans, or if there is tuned/pulsed beam in the upstream beamline.
    • ON after an initial beam tune has been established and if there are only minor steering adjustments and “tweaks” being made.
    • If in doubt, call FTOF expert

FTT (mmft1) recovery:

    • power cycle FTT LV (FT->FTT->turn LV OFF, turn LV ON)
    • RunControl "reset"
    • roc_reboot mmft1

Torus and/or Solenoid Fast Dump

    1. Notify MCC to request beam OFF and to drop Hall B status to Power Permit
    2. Call Engineering on-call
    3. Make separate log entry with copies to HBTORUS and HBSOLENOID logbooks. In the "Notify" field add Ruben Fair, Probir Goshal, Dave Kashy and
    4. Notify Run Coordinator
    5. Turn off all detectors

Beam to the Faraday Cup: Fast Shut Down elements

  • Upstream, Midstream, Downstream, BOM, and Solenoid FSD elements should always be in the state UNMASKED
  • No changes to the FSD threshold should be made without RC or beamline expert approval

Turning MVT off: do not press "All HV OFF". Instead go to 'Restore settings' from the MVT overview screen:

    1. SafeMode.snp for beam tuning and Moeller runs
    2. MV_HV_FullField.snp for full solenoid field
    3. MV_HV_MidField.snp wgeb solenoid < 4T

CND CAMAC crate switch on / reboot:

If CAMAC crate (camac1) holding CND CFD boards is switched off for any reason, always reset the associated CND CFD thresholds by typing, ssh into clondaq3 and type:

$CODA/src/rol/Linux_x86_64/bin/cnd-cfd-thresh -w 0

If this command is failing and the crate is not responding, first reboot it as follows: roc_reboot camac1

and then repeat the threshold reset command above.

Solutions for common SVT issues: see [9]

RICH recovery procedures: [10]

To be followed in case of:

  1. DAQ crash: rich4 is not responding or
  2. RICH alarms (LV,missing tile, temperature, etc).

If it does not work or you are uncertain about what to do, contact the RICH expert on call. Please note that missing tiles typically occur due to lost communication. Keep in mind that the recovery procedure will kill DAQ. If DAQ is running for purpose other than data taking (for which the RICH acceptance is important) do not initiate the recovery procedure. The most critical parameter for the RICH is the temperature of the photosensors. If the temperature rises above limits, an interlock will automatically turn the RICH HV and LV off. If this happens, notify the expert on call and keep taking data without RICH.

Observed asymmetry in FTCal ADC Scalers

When running at low current (<10 nA), FTC asymmetries can be affected by hot channels. If large asymmetries are detected in those conditions, cross check with other detectors (BOM rates, DC occupancies) before taking further action.

CED display not showing events

Sometimes, CED screens are up but not active (events not showing at all). To solve this:

  1. Close the CED sceen
  2. ssh to clondaq4
  3. Once in the right machine, type ced, and wait for the windows to appear.
  4. Go to "File" -> "Connect to ET ring". Make sure to select clondaq6, as the online data is being streamed.
  5. Go to "Events" (same level as "File") and check "Auto Next-Event Every 2.0 sec" (if it is not selected).

You should see the events now, if all is working well.